OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by this equation: y=-2x^2-4x+6 Write your answer as two ordered pairs: (x1,y1),(x2,y2) Separate the values with a comma. Round, if necessary, to the nearest hundredth. @lucaz

OpenStudy (lucaz):

factor out -2 -2(x^2+2x-3) an easy way to factor quadratics is this Ax^2+Bx+C=0, where A=1, B=2 and C=-3 now find two numbers a and b that the product is A*C and the sum is B, then the factored form will be (x+a)(x+b) A*C=1*-3=-3 and B=2 try the numbers 3 and -1 3*-1=-3 and 3+(-1)=2, they work, so the factored form is (x+3)(x-1). then you have -2((x+3)(x-1))=0 x+3=0 -> x=-3 x-1=0 -> x=1 the x-intercepts are (-3,0) and (1,0) but there are some quadratic equations that cannot be factored, then you will have to use the quadractic formula: